OpenStudy (anonymous):

y/3 minus 9 =12 what is y?

OpenStudy (anonymous):

y/3 - 9 = 12 add 9 both sides then divide both sides by 3

OpenStudy (studygurl14):

\[\frac{y}{3}-9=12\] \[y - 9(3) = 12(3) \rightarrow y-27=36\rightarrow y = 63\]

OpenStudy (anonymous):

y=63

OpenStudy (studygurl14):

you can do it a couple different ways

OpenStudy (studygurl14):

@sangya21 's is one way, mine is another

OpenStudy (anonymous):

y=63

OpenStudy (studygurl14):

i think you mean multiply b 3 @sangya21 , not divide...?

OpenStudy (anonymous):

@StudyGurl14 thanks for showing me. Yups its multiply.

OpenStudy (studygurl14):

:)